OpenStudy (anonymous):

find the area of the region in the first quadrant that is bounded above by the graph of f(x) = sqrt(x) and below by the x-axis and the line g(x) = x - 2 using integration with respect to y. show steps/explain.

OpenStudy (turingtest):

draw it, always find the points of intersection once you have all that it is easier to approach the problem

OpenStudy (anonymous):

how do i find y upper and lower

OpenStudy (anonymous):

ya, i've got a rough sketch like that

OpenStudy (anonymous):

looks like 2 and 0

OpenStudy (anonymous):

for the y-axis

OpenStudy (anonymous):

well... u have an exact graph.. how do i find those values otherwise

OpenStudy (anonymous):

find the intercsection

OpenStudy (turingtest):

set the two formulas equal

OpenStudy (anonymous):

and find both coordinates

OpenStudy (turingtest):

\[y^2=y+2\implies y^2-y-2=(y+1)(y-2)=0\]and we don't care about the negative intersect since we're in QI

OpenStudy (turingtest):

I solved it for y because it's easier than solving \[\sqrt x=x-2\], and also because we are looking for the y-value anyway

OpenStudy (anonymous):

how'd you get y^2 = y + 2

OpenStudy (anonymous):

oh... i think i got it

OpenStudy (turingtest):

\[\sqrt x=y\implies x=y^2\]\[x-2=y\implies y+2=x\]setting\[x=x\]we get\[y^2=y+2\]

OpenStudy (turingtest):

now how do we decide which is our upper and which is our lower do you know?

OpenStudy (anonymous):

0 lower, 2 upper... cause 2 is bigger than 0 and it looks right?

OpenStudy (turingtest):

yes, but which function minus which function in the integrand?

OpenStudy (anonymous):

right most - left most. so here it's y+2 - y^2

OpenStudy (turingtest):

yep

OpenStudy (turingtest):

well that's all happy integrating!

OpenStudy (anonymous):

how come i dont have to split it into two separate pieces like in the "with respect to x" version

OpenStudy (turingtest):

because I neglected to mention that part :P

OpenStudy (anonymous):

lol

OpenStudy (turingtest):

oh, no, it;s because there is no "chink" in the "right-left" like there is in the "top-bottom" way along x

OpenStudy (anonymous):

it might be right.. i was just wondering

OpenStudy (anonymous):

define chink, lol

OpenStudy (turingtest):

sharp corner-type pointy thing ^^^ oxford english standard

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

meh, well i guess i'll just have to eye-ball it

OpenStudy (turingtest):

very rough sketch as we gio along x we are adding up little rectangles that stretch up and down

OpenStudy (anonymous):

... eye-balling it probably not good for bridge building :)

OpenStudy (anonymous):

yeah, i gotcha

OpenStudy (turingtest):

but when we get to the point we have to use a different bottom....